OpenStudy (brittany):

In two years, Linda will be half as old as her mother. Her mother is 38 years old now. What is Linda's current age?

OpenStudy (anonymous):

Linda is 17.

OpenStudy (anonymous):

x+2=(38+2)/2 x+2=20 x=18